Can someone help with parts c-e?A 3000 kg rhinoceros running north at 10 m/s collides head on with an 800-kg Land Rover moving east at 25 m/s. The two stick
Answer #1

c. We know that Momentum remains conserve in head on collision.

So, (M​​V)Rhinoceros + (MV)Land rover = (MV)Combined

=> (3000 kg)*(10 m/s j) + (800 kg)*(25 m/s i) = (3000+800) kg * V​​​​​​Combined

​​​​​​=> V​​​​​Combined = (7.9 j + 5.3 i) kg-m/s

Hence, Combined Velocity = (5.3i + 7.9j) (kg-m)/s

d. Combined Speed = √[(5.3)² + (7.9)²] m/s = 9.49 m/s

Direction of system = (5.3i + 7.9j)/9.49

e. KE​​​​​Before = (KE)Rhinoceros + (KE)Land rover = (0.5)[(3000 kg)*(10 m/s)² + (800 kg)*(25 m/s)²] = 265000 J

KE​​​​​After = 0.5*((3000+8000)kg * (9.49 m/s)²) = 171114 J